# A single-loop SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization

### v1

@@ 1.3 Contributions
1.4 Notations and definitions
@@
(7) $\phi^{\prime}({\mathbf{x}};{\mathbf{v}})\geq{\mathbf{v}}^{\top}{\boldsymbol{\gamma}},\text{ for any }{\boldsymbol{\gamma}}\in\partial\phi({\mathbf{x}}).$
Definition 1.1 (Moreau Envelope [13]).
For a $\rho_{\phi}$-weakly convex function $\phi$, the Moreau envelope and the proximal mapping for any $t\in(0,\frac{1}{\rho_{\phi}})$ are defined by $\phi_{t}({{\mathbf{x}}}):=\min_{{\mathbf{z}}}\left\{\phi({{\mathbf{z}}})+\frac{1}{2t}\|{{\mathbf{z}}}-{{\mathbf{x}}}\|^{2}\right\}$ and $\operatorname{Prox}_{t\phi}({{\mathbf{x}}}):=\arg\min_{\mathbf{z}}\{\phi({{\mathbf{z}}})+\frac{1}{2t}\|{{\mathbf{z}}}-{{\mathbf{x}}}\|^{2}\}.$
We give the definition of stationarity as follows. In deterministic cases—where randomness is not involved—the expectation operator $\mathbb{E}$ should be removed from the definitions below.
Definition 1.2 (Stationarity).
Let $\epsilon>0$ and $\delta>0$, and suppose $\psi$ and $\phi$ are $\rho_{\psi}$- and $\rho_{\phi}$-weakly convex functions.
$\mathrm{(a)}$ We call ${\mathbf{x}}^{*}$ a near-$\epsilon$ stationary point of $\min_{{\mathbf{x}}}\phi({\mathbf{x}})$ in expectation, if ${\mathbf{x}}^{*}\in{\mathrm{dom}}(\phi)$, and there exists $0<t<\frac{1}{\rho_{\phi}}$, such that $\mathbb{E}[\|\nabla\phi_{t}({\mathbf{x}}^{*})\|^{2}]=\mathbb{E}[\|\frac{1}{t}(\operatorname{Prox}_{t\phi}({{\mathbf{x}}^{*}})-{\mathbf{x}}^{*})\|^{2}]\leq~{}\epsilon^{2}.$
$\mathrm{(b)}$ We call ${\mathbf{x}}^{*}$ an $\epsilon$-KKT point of problem $\min_{{\mathbf{x}}:\psi({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation, if it holds that ${\mathbf{x}}^{*}\in{\mathrm{dom}}(\phi)\cap{\mathrm{dom}}(\psi)$, and there exists $\lambda_{{\mathbf{x}}^{*}}\geq 0$ such that
@@
$\mathbb{E}\left[\textnormal{dist}(\mathbf{0},\partial\phi({\mathbf{x}}^{*})+\lambda_{{\mathbf{x}}^{*}}\partial\psi({\mathbf{x}}^{*}))\right]\leq\epsilon,\,\,\mathbb{E}\left[\psi({\mathbf{x}}^{*})\right]\leq\epsilon,\text{ and }|\lambda_{{\mathbf{x}}^{*}}\psi({\mathbf{x}}^{*})|\leq\epsilon.$
$\mathrm{(c)}$ We call ${\mathbf{x}}^{*}$ a $(\delta,\epsilon)$-KKT point of problem $\min_{{\mathbf{x}}:\psi({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation, if it holds that ${\mathbf{x}}^{*}\in{\mathrm{dom}}(\phi)\cap{\mathrm{dom}}(\psi)$, and $\mathbb{E}[\|{\mathbf{x}}^{*}-\widetilde{{\mathbf{x}}}^{*}\|]\leq\delta$ with $\widetilde{{\mathbf{x}}}^{*}$ being an $\epsilon$-KKT point of problem $\min_{{\mathbf{x}}:\psi({\mathbf{x}})\leq 0}\phi({\mathbf{x}})$ in expectation.
1.5 Organizations
The rest of this paper is organized as follows. Section 2 introduces the exact penalty models. In Section 3, we propose a single-loop SPIDER-type stochastic subgradient method and establish its iteration complexity results. Numerical experiments are presented in Section 4, showcasing the performance of our method. Finally, Section 5 concludes the paper.
2 Model analysis
In this section, we establish the relationships among models Eq. PEq. 5 and Eq. PP, which are presented in the following diagram for the convenience of the reader. Details are given in the mentioned lemmas and theorem.

### v2

@@ 2 Model analysis
2.1 Regularity conditions
To derive the exact penalization result, we assume the following (uniform) Slater-type CQ, which holds for tested instances of Examples 1.1 and 1.2 in Section 4.
Assumption 2 ((uniform) Slater-type CQ [27, 46]).
There exist $B_{g}>~0$, $B>0$, and $\overline{\rho}>\rho_{g}\geq 0$, such that for any ${\mathbf{x}}{\in{\mathcal{X}}}$ satisfying $0<g_{+}({\mathbf{x}})\leq B_{g}$ with $g_{+}$ defined in (6), it holds
(8) $\max_{i\in[m]}g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B,\text{ for some }\mathbf{y}\in\mathrm{relint}(\mathcal{X}).$
Below we give the relationships among the Slater-type CQ, Slater’s condition and the strong feasibility condition assumed in [10], where the latter two conditions are stated as follows:
(9) $\text{Slater's condition: }\exists\,{\mathbf{y}}_{\mathrm{feas}}\in\mathrm{relint}({\mathcal{X}})\text{ such that }\max_{i\in[m]}g_{i}({\mathbf{y}}_{\mathrm{feas}})<0;$
(10) $\text{strong feasibility: }\exists\,{\mathbf{y}}_{\mathrm{feas}}\in\mathcal{X}\text{ such that }\max_{i\in[m]}g_{i}({\mathbf{y}}_{\mathrm{feas}})\leq-2\rho_{g}D^{2}.$
Proposition 2.1.
Under 1, the following statements hold:
  • $(\mathrm{i})$
    If $\{g_{i}\}_{i\in[m]}$ are convex, Slater’s condition and the Slater-type CQ are equivalent.
  • $(\mathrm{ii})$
    If the strong feasibility condition holds, so does the Slater-type CQ.
Proof 2.2.
(i) Suppose Slater’s condition holds. Let ${\mathbf{y}}={\mathbf{y}}_{\mathrm{feas}},$ $\overline{\rho}=\frac{-\max_{i\in[m]}g_{i}({\mathbf{y}}_{\mathrm{feas}})}{D^{2}},$ $B=\frac{-\max_{i\in[m]}g_{i}({\mathbf{y}}_{\mathrm{feas}})}{2}$, and $B_{g}$ be any positive constant. Then Eq. 8 trivially holds with the chosen $\mathbf{y}$}, $\overline{\rho}$, $B$, and any ${\mathbf{x}}\in{\mathcal{X}}$. Thus Slater-type CQ holds since $\rho_{g}=0<\overline{\rho}$.
Conversely, if the Slater-type CQ holds, then by definition there exist ${\mathbf{y}}\in\mathrm{relint}({\mathcal{X}})$ and $B>0$ such that $\max_{i\in[m]}g_{i}({\mathbf{y}})\leq-B<0,$ which immediately implies Slater’s condition. Hence, we establish the equivalence.
(ii) Because ${\mathcal{X}}\subset\mathbb{R}^{d}$ is nonempty convex and ${\mathbf{y}}_{\mathrm{feas}}\in\mathcal{X}$, there must exist ${\mathbf{y}}\in\mathrm{relint}({\mathcal{X}})$ such that $\|{\mathbf{y}}-{\mathbf{y}}_{\mathrm{feas}}\|\leq\frac{\rho_{g}D^{2}}{2l_{g}}$. By the $l_{g}$-Lipschitz continuity of each $g_{i}$ from 1(c), it holds $g_{i}(\mathbf{y})-g_{i}(\mathbf{y}_{\mathrm{feas}})\leq l_{g}\|{\mathbf{y}}-{\mathbf{y}}_{\mathrm{feas}}\|\leq\rho_{g}D^{2}/2,\,\forall\,i\in[m]$. Thus $\max_{i\in[m]}g_{i}(\mathbf{y})\leq\max_{i\in[m]}g_{i}(\mathbf{y}_{\mathrm{feas}})+\rho_{g}D^{2}/2$. Now let $\overline{\rho}=2\rho_{g}$. We have for any ${\mathbf{x}}\in{\mathcal{X}}$ that
$\max_{i\in[m]}g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$ $\leq\max_{i\in[m]}g_{i}(\mathbf{y}_{\mathrm{feas}})+\frac{\rho_{g}D^{2}}{2}+\rho_{g}\|\mathbf{y}-\mathbf{x}\|^{2}$
$\leq-2\rho_{g}D^{2}+\frac{\rho_{g}D^{2}}{2}+\rho_{g}D^{2}=-\frac{\rho_{g}D^{2}}{2}.$
Hence, the Slater-type CQ holds with $B=\frac{\rho_{g}D^{2}}{2}$, $\overline{\rho}=2\rho_{g}$, and any $B_{g}>0$.
We highlight that the strong feasibility condition in Eq. 10 ensures the Slater-type CQ with any $B_{g}>0$. Later in the proof of Theorem 2.9, we assume Slater-type CQ with $B_{g}>m\rho_{g}D^{2}$, which is still implied by the strong feasibility condition. However, the converse does not hold in general. The following example illustrates that a Slater-type CQ with $B_{g}>m\rho_{g}D^{2}$ does not imply the strong feasibility condition.
Example 2.3.
Consider the nonconvex function $g(x)=-6+8x+x^{2}-x^{3}$ over the set $\mathcal{X}=[0,1].$ Since $g^{\prime\prime}(x)=2-6x$, it holds $g^{\prime\prime}(x)\geq-4,\,\forall x\in{\mathcal{X}}$. Then $g$ is $4$-weakly convex on $\mathcal{X}$}. Take $y=0.01\in\mathrm{relint}({\mathcal{X}})$. It is straightforward to verify that $g(y)+\frac{5}{2}|x-y|^{2}\leq-3$ for any $x\in\mathcal{X}$. Thus, the Slater-type CQ holds with parameters $B_{g}=1000,B=3$, and $\overline{\rho}=5>4=\rho_{g}.$ Thus, we have $B_{g}>\rho_{g}mD^{2}$, where $m=1$, and $D=1$. However, the strong feasibility condition fails because it holds that $\min_{x\in[0,1]}g(x)=-6\;>\;-8=-2\rho_{g}D^{2}.$
2.2 Exact penalization with penalty function $\cdot$$_{+}$
In this subsection, we study the relationship between Eq. P and Eq. 6. Since each $g_{i}$ is $\rho_{g}$-weakly convex for $i\in[m]$, and $f$ is $\rho_{f}$-weakly convex, $F$ in Eq. 6 is $\rho$-weakly convex with $\rho:=\rho_{f}+\beta m\rho_{g}$.
First, we present a key lemma as follows, whose proof is given in Appendix A.
Lemma 2.4.
Under 1 and 2, let $\theta=\sqrt{2B(\overline{\rho}-\rho_{g})}$ and ${\mathbf{x}}\in\mathcal{X}$ satisfy $0<g_{+}({\mathbf{x}})\leq B_{g}$. Given $\mathcal{J}_{1}\subset[m]$ such that $\mathcal{J}_{1}\cap\{i:g_{i}({\mathbf{x}})\geq 0\}\neq\emptyset$, let $\mathcal{J}_{2}=[m]\setminus\mathcal{J}_{1}$ and choose $\{\lambda_{i}\}_{i\in\mathcal{J}_{2}}\subset\mathbb{R}_{+}$ satisfying $\lambda_{i}g_{i}({\mathbf{x}})\geq 0$ for all $i\in\mathcal{J}_{2}$. Then,
$\min_{\{\boldsymbol{\zeta}_{g_{i}}\in\partial[g_{i}({\mathbf{x}})]_{+}\}_{i\in\mathcal{J}_{1}},\{\boldsymbol{\zeta}_{g_{i}}\in\partial g_{i}({\mathbf{x}})\}_{i\in\mathcal{J}_{2}},{\mathbf{u}}\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}})}\left\|\sum_{i\in\mathcal{J}_{1}}\zeta_{g_{i}}+\sum_{i\in\mathcal{J}_{2}}\lambda_{i}\zeta_{g_{i}}+{\mathbf{u}}\right\|\geq\theta.$
Now, we establish the relationship between the $\epsilon$-stationary point of problem Eq. 6 and the $\epsilon$-KKT point of Eq. P.
Lemma 2.5.
Suppose 1 and 2 hold. Given $\epsilon>0$, let the penalty parameter $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$ in Eq. 6 with $\theta=\sqrt{2B(\overline{\rho}-\rho_{g})}$. If ${\mathbf{x}}^{*}\in{\mathcal{X}}$ satisfying $g_{+}({\mathbf{x}}^{*})\leq B_{g}$ is an $\epsilon$-stationary point of problem Eq. 6, then $\mathbf{x}$$^{*}$ satisfies $g_{i}({\mathbf{x}}^{*})\leq 0$ for all $i\in[m]$, and is an $\epsilon$-KKT point of problem Eq. P.
Proof 2.6.
Since $\mathbf{x}$$^{*}$ is an $\epsilon$-stationary point of problem Eq. 6, there exist ${\boldsymbol{\gamma}}\in\partial f({\mathbf{x}}^{*})$, $\{{\boldsymbol{\zeta}}_{g_{i}}\in\partial[g_{i}({\mathbf{x}}^{*})]_{+}\}_{i\in[m]}$, and ${\mathbf{u}}\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{*})$ such that
(11) $\left\|\textstyle\beta\sum_{i=1}^{m}{\boldsymbol{\zeta}}_{g_{i}}+{\mathbf{u}}+{\boldsymbol{\gamma}}\right\|\leq\epsilon.$
We claim $g_{i}({\mathbf{x}}^{*})\leq 0$ for all $i\in[m]$. Otherwise, it holds $0<g_{+}({\mathbf{x}}^{*})\leq B_{g}$ and $m]\cap\{i:g_{i}({\mathbf{x}}^{*})\geq 0\}\neq\emptyset$. We then derive from Eq. 11 and the triangle inequality that
(12) $\epsilon$ $\geq{\beta}\left\|\textstyle\sum_{i=1}^{m}{\boldsymbol{\zeta}}_{g_{i}}+\frac{{\mathbf{u}}}{\beta}\right\|-\|{\boldsymbol{\gamma}}\|\geq{\beta}\left\|\textstyle\sum_{i=1}^{m}{\boldsymbol{\zeta}}_{g_{i}}+\frac{{\mathbf{u}}}{\beta}\right\|-l_{f}\geq\beta\theta-l_{f},$
where the second inequality holds by 1(c), and the third one follows from Lemma 2.4 with ${\mathcal{J}}_{1}=[m]$. This contradicts the condition $\beta>\frac{1}{\theta}(l_{f}+\epsilon)$. Hence $g_{i}({\mathbf{x}}^{*})\leq 0$ for all $i\in[m]$.
Now we find a corresponding multiplier as follows. For each $i\in[m]$, if $g_{i}({\mathbf{x}}^{*})=0$, we obtain $\partial[g_{i}({\mathbf{x}}^{*})]_{+}={\mathrm{co}}(\partial g_{i}({\mathbf{x}}^{*})\cup\{\mathbf{0}\})$ by [18, Proposition 2.3.12] and the weak convexity of $g_{i}]_{+}$. Thus, there exists $\lambda_{i}\in[0,1]$ such that ${\boldsymbol{\zeta}}_{g_{i}}\in\lambda_{i}\partial g_{i}({\mathbf{x}}^{*})$. In this case, we set $t_{i}=\lambda_{i}$. If $g_{i}({\mathbf{x}}^{*})<0$, we have $\partial[g_{i}({\mathbf{x}}^{*})]_{+}=\{\mathbf{0}\}$ so ${\boldsymbol{\zeta}}_{g_{i}}=\mathbf{0}$, and we set $t_{i}=0$. For the constructed $\mathbf{t}$}, we have $t_{i}g_{i}({\mathbf{x}}^{*})=0,\forall\,i\in[m]$. In addition, $\beta\sum_{i=1}^{m}{\boldsymbol{\zeta}}_{g_{i}}\in\sum_{i=1}^{m}\beta t_{i}\partial g_{i}({\mathbf{x}}^{*})$, and thus Eq. 11 indicates $\textnormal{dist}\big{(}\mathbf{0},\partial f({\mathbf{x}}^{*})+\sum_{i=1}^{m}\beta t_{i}\partial g_{i}({\mathbf{x}}^{*})+{\mathcal{N}}_{\mathcal{X}}({\mathbf{x}}^{*})\big{)}\leq\epsilon$. Therefore $\mathbf{x}$$^{*}$ is an $\epsilon$-KKT point of problem Eq. P with the multiplier $\beta{\mathbf{t}}$.
2.3 Exact penalization with a “smoothed” penalty approach
In this subsection, we establish the relationship between models Eq. P and Eq. PP. We begin by showing that the function $F^{\nu}$ in Eq. PP is weakly convex.
Lemma 2.7.
Suppose 1 holds. For any $\nu>0$, the function $F^{\nu}$ defined in Eq. PP is $\rho$-weakly convex, where $\rho=\rho_{f}+\beta m\rho_{g}$.
Proof 2.8.
To obtain the desired result, we only need to prove $h^{\nu}({\mathbf{g}}({\mathbf{x}}))$ is $m\rho_{g}$-weakly convex. Notice that
$h^{\nu}({\mathbf{g}}({\mathbf{x}}))+\frac{m\rho_{g}}{2}\|{\mathbf{x}}\|^{2}=\sum_{i=1}^{m}\max_{0\leq y\leq 1}\left[y\cdot g_{i}({\mathbf{x}})-\frac{\nu}{2}y^{2}+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}\right].$
Since $g_{i}$ is $\rho_{g}$-weakly convex for each $i\in[m]$, $y\cdot g_{i}({\mathbf{x}})-\frac{\nu}{2}y^{2}+\frac{\rho_{g}}{2}\|{\mathbf{x}}\|^{2}$ is convex with respect to $\mathbf{x}$} for all $0\leq y\leq 1$. Thus by Danskin’s theorem [9], $h^{\nu}({\mathbf{g}}({\mathbf{x}}))+\frac{m\rho_{g}}{2}\|{\mathbf{x}}\|^{2}$ is convex.
Next, we introduce some notations and key properties of the Moreau envelope. Denote $F_{\frac{1}{2\rho}}$ and $F^{\nu}_{\frac{1}{2\rho}}$ as the Moreau envelope of $F$ and $F^{\nu}$ with parameter $\frac{1}{2\rho}$ on $\mathcal{X}$, respectively, where $\rho=\rho_{f}+\beta m\rho_{g}$. We let
(13) $\overline{{\mathbf{x}}}^{*}=\arg\min_{{\mathbf{x}}\in\mathcal{X}}F({\mathbf{x}})+\rho\|{\mathbf{x}}-{\mathbf{x}}^{*}\|^{2},\quad\widehat{{\mathbf{x}}}^{*}=\arg\min_{{\mathbf{x}}\in\mathcal{X}}F^{\nu}({\mathbf{x}})+\rho\|{\mathbf{x}}-{\mathbf{x}}^{*}\|^{2}.$
Then we have
(14) $\nabla F_{\frac{1}{2\rho}}({\mathbf{x}}^{*})=2\rho({\mathbf{x}}^{*}-\overline{{\mathbf{x}}}^{*}),\ \nabla F^{\nu}_{\frac{1}{2\rho}}({\mathbf{x}}^{*})=2\rho({\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}^{*}),\text{ and}$
(15) $\mathbf{0}\in\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\beta\sum_{i=1}^{m}\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}\left(\widehat{\mathbf{x}}^{*}\right)}{\nu}\right)\partial g_{i}\left(\widehat{\mathbf{x}}^{*}\right)+2\rho\left(\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\right)+\mathcal{N}_{\mathcal{X}}(\widehat{\mathbf{x}}^{*}).$
In addition, by [21, Eqn. (4.3)], it holds that
@@ 2.3 Exact penalization with a “smoothed” penalty approach
We now present our main results to demonstrate the effectiveness of solving a smoothed exact penalty problem for both deterministic and stochastic settings.
Theorem 2.9.
Suppose that 1 holds and 2 holds with $B_{g}>m\rho_{g}D^{2}$. Given $\epsilon>0$, let $\theta=\sqrt{2B(\overline{\rho}-\rho_{g})}$, $\rho=\rho_{f}+\beta m\rho_{g}$,
(17) $\beta>\max\left\{\frac{1}{\theta}(l_{f}+\epsilon),\frac{2(B_{f}+\rho_{f}D^{2})}{B_{g}-m\rho_{g}D^{2}}\right\},0<\overline{\epsilon}\leq\epsilon\min\left\{\frac{\beta\theta-l_{f}}{2B_{g}},\frac{\beta\theta-l_{f}}{m\beta l_{g}+\beta\theta-l_{f}},2\rho\right\},$
$\text{ and }0<\nu\leq\min\left\{B_{g},\frac{\epsilon}{\beta},\frac{B_{g}-m\rho_{g}D^{2}}{2m},\frac{\epsilon}{2}\right\}.$
If $\mathbf{x}$$^{*}$ is a near-$\overline{\epsilon}$ stationary point of problem Eq. PP with $t=\frac{1}{2\rho}=\frac{1}{2(\rho_{f}+\beta m\rho_{g})}$ deterministically or in expectation, then $\|{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}^{*}\|\leq\epsilon$ holds deterministically or in expectation, where $\widehat{{\mathbf{x}}}^{*}$ is defined in Eq. 13. In addition, $\widehat{{\mathbf{x}}}^{*}$ is an $\epsilon$-KKT point of problem Eq. P deterministically or in expectation.
Proof 2.10.
We only need to prove the theorem in the expectation case, when $\mathbf{x}$$^{*}$ is a random vector. For the deterministic case, one can simply drop the expectation notation $\mathbb{E}$.
First, from Definition 1.3 and Eq. 14, it follows $\mathbb{E}\left[\|2\rho(\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*})\|^{2}\right]\leq~\overline{\epsilon}^{2}$ by the near-$\overline{\epsilon}$ stationarity of $\mathbf{x}$$^{*}$ in expectation. Because $\overline{\epsilon}\leq 2\rho\epsilon$, it follows that
Second, we define two index sets: $\mathcal{J}_{2}=\{i:g_{i}(\widehat{\mathbf{x}}^{*})/\nu\leq 1\}$ and $\mathcal{J}_{1}=[m]\setminus\mathcal{J}_{2}$. Construct the vector $\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}}$ by setting $(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}=\beta\mathrm{Proj}_{[0,1]}(g_{i}(\widehat{\mathbf{x}}^{*})/\nu)\geq 0$ for $i\in\mathcal{J}_{2}$ and $(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}=0$ for $i\in\mathcal{J}_{1}$. We next bound $|(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})|$ for each $i$. For $i\in\mathcal{J}_{1}$, we directly have $(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})=0$. For $i\in\mathcal{J}_{2}$, if $g_{i}(\widehat{\mathbf{x}}^{*})\leq 0$, then $(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}=0$ and $(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})=0$; if $g_{i}(\widehat{\mathbf{x}}^{*})>0$, we have from $g_{i}(\widehat{\mathbf{x}}^{*})/\nu\leq 1$ that $g_{i}(\widehat{\mathbf{x}}^{*})\leq\nu\leq\epsilon/2$, and thus $|(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})|=(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})\leq\beta g_{i}(\widehat{\mathbf{x}}^{*})\leq\beta\nu\leq\epsilon.$ Hence,
(19) $|(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})|\leq\epsilon,\forall\,i\in[m].$
Third, we show $g_{+}(\widehat{{\mathbf{x}}}^{*})\leq B_{g}$. By Eq. 15 and the definition of ${\mathcal{J}}_{1}$ and ${\mathcal{J}}_{2}$, we have
(20) $\mathbf{0}\in\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\beta\partial\sum_{i\in\mathcal{J}_{1}}[g_{i}\left(\widehat{\mathbf{x}}^{*}\right)]_{+}+\sum_{i\in\mathcal{J}_{2}}(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}\partial g_{i}\left(\widehat{\mathbf{x}}^{*}\right)+\mathcal{N}_{\mathcal{X}}(\widehat{\mathbf{x}}^{*})+2\rho\left(\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\right).$
By 1, there exists $\mathbf{x}_{\mathrm{feas}}\in{\mathcal{X}}\cap\{{\mathbf{x}}:{\mathbf{g}}({\mathbf{x}})\leq 0\}$. Hence, from Eq. 13, it holds
(21) $f\left(\widehat{\mathbf{x}}^{*}\right)+\beta h^{\nu}\left({\mathbf{g}}\left(\widehat{\mathbf{x}}^{*}\right)\right)+\rho\left\|\widehat{\mathbf{x}}^{*}-\mathbf{x}^{*}\right\|^{2}$
$\leq$ $f\left(\mathbf{x}_{\mathrm{feas}}\right)+\beta h^{\nu}\left({\mathbf{g}}\left(\mathbf{x}_{\mathrm{feas}}\right)\right)+\rho\left\|\mathbf{x}_{\mathrm{feas}}-\mathbf{x}^{*}\right\|^{2}=f\left(\mathbf{x}_{\mathrm{feas}}\right)+\rho\left\|\mathbf{x}_{\mathrm{feas}}-\mathbf{x}^{*}\right\|^{2}.$
Thus, $f(\widehat{{\mathbf{x}}}^{*})+\beta h^{\nu}({\mathbf{g}}(\widehat{{\mathbf{x}}}^{*}))+\rho\|\widehat{{\mathbf{x}}}^{*}-{\mathbf{x}}^{*}\|^{2}\leq f({\mathbf{x}}_{\mathrm{feas}})+\rho D^{2}$ by 1(a). In addition, we have $\beta\geq\frac{B_{f}+\rho_{f}D^{2}}{B_{g}-m\nu-m\rho_{g}D^{2}}$ by $\beta>\frac{2(B_{f}+\rho_{f}D^{2})}{B_{g}-m\rho_{g}D^{2}}$ and $\nu\leq\frac{B_{g}-m\rho_{g}D^{2}}{2m}$. Now further by $m\rho_{g}D^{2}<B_{g}$, $\rho=\rho_{f}+\beta m\rho_{g}$, and Eq. 1, it follows that
(22) $\textstyle h^{\nu}({\mathbf{g}}(\widehat{{\mathbf{x}}}^{*}))\leq\frac{1}{\beta}\Bigl{(}f({\mathbf{x}}_{\mathrm{feas}})-f(\widehat{{\mathbf{x}}}^{*})+\rho D^{2}\Bigr{)}\leq\frac{B_{f}+\rho D^{2}}{\beta}\leq B_{g}-m\nu.$
This implies $g_{+}(\widehat{{\mathbf{x}}}^{*})\leq h^{\nu}({\mathbf{g}}(\widehat{{\mathbf{x}}}^{*}))+m\nu\leq B_{g}$ by the definition of $h^{\nu}$.
We now discuss two events: ${\mathcal{A}}_{1}$ as the event of $\mathcal{J}_{1}=\emptyset$ and ${\mathcal{A}}_{2}$ as the event of $\mathcal{J}_{1}\neq\emptyset$. Let $\mathrm{Prob}_{1}$ and $\mathrm{Prob}_{2}$ denote their probabilities. When ${\mathcal{A}}_{1}$ happens, we have $\textnormal{dist}\left(\mathbf{0},\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\sum_{i=1}^{m}(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}\partial g_{i}\left(\widehat{\mathbf{x}}^{*}\right)+\mathcal{N}_{\mathcal{X}}(\widehat{\mathbf{x}}^{*})\right)\leq 2\rho\|\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|$ by Eq. 20. Hence,
@@
(23) $\textnormal{dist}\left(\mathbf{0},\partial f\left(\widehat{\mathbf{x}}^{*}\right)+\sum_{i=1}^{m}(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}\partial g_{i}\left(\widehat{\mathbf{x}}^{*}\right)+\mathcal{N}_{\mathcal{X}}(\widehat{\mathbf{x}}^{*})\right)\leq 2\rho\|\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|,$
$g_{i}(\widehat{\mathbf{x}}^{*})\leq\frac{\epsilon}{2},\text{ and }|(\boldsymbol{\lambda}_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}\left(\widehat{\mathbf{x}}^{*}\right)|\leq\epsilon,\forall\,i\in[m].$
When ${\mathcal{A}}_{2}$ happens, we have $\mathcal{J}_{1}\cap\{i:g_{i}(\widehat{\mathbf{x}}^{*})\geq 0\}\neq\emptyset$, and $(\lambda_{\widehat{\mathbf{x}}^{*}})_{i}g_{i}(\widehat{\mathbf{x}}^{*})\geq 0$ for all $i\in\mathcal{J}_{2}$. Using Lemma 2.4 and 1(c), we obtain that
(24) $2\rho\|\widehat{\mathbf{x}}^{*}-{\mathbf{x}}^{*}\|$